a car traveling at 70 miles per hour, what is the distance of travel within 30 seconds?
Found 2 solutions by thejackal, DrBeeee:
Answer by thejackal(72)

convert 30 seconds to hours = 0.00833333
Distance traveled = time * speed = 0.00833333 * 70 = 0.583331 miles

Answer by DrBeeee(684)

Let r = speed or rate in mph
Let t = time travelled in hrs
Let d = distance travelled in t hours, in miles
Now use the distance formula
(1) d = rt
The rate is 70 miles per hour, but the time is 30 seconds. We need the time in hours, not seconds, so how do you change 30 sec to x hrs?
Do this
(2) 30 sec = 30 sec * x hr/sec
Since one hour is for 60 sec/min*60 min/hr or 3600 sec/hr, we have the inverse or
(3) x = 1/3600 hr/sec
Then we use (1) to get
(4) d = 70*(30/3600) or
(5) d = 7/12
Answer: the car travels 7/12 mile in 30 seconds. This is 3080 feet. Note that in one second at 70 mph you travel about 103 feet.